OpenStudy (anonymous):

7) A football is booted with an initial velocity of 18 m/s 55o above the horizontal. a. What is its speed at the top of its arc? 10.3 m/s b. How far away will the football land?

OpenStudy (anonymous):

I already have part a done. Im just stuck on part b.

OpenStudy (anonymous):

Find the total time the ball is in the air. Multiply the time by the *horizontal* velocity of the ball.

OpenStudy (anonymous):

How do I find the time?

OpenStudy (anonymous):

Use v = v_0 -g*t

OpenStudy (anonymous):

so V = 14.74 - 9.8t 0 = 14.74 - 9.8t t = 1.5 d = 30.9

OpenStudy (anonymous):

You are correct. Thank you. You were the only one who posted anything.

OpenStudy (anonymous):

1.5 seconds is the time when the football is at the top (or halfway) through the arc. You need to double the time to get the total time in the air.

OpenStudy (anonymous):

Yeah, I forgot to show that step in my work, but I understand the concept and got the right answer. Thanks!

OpenStudy (anonymous):

Cool. No problem.